how do i solve this system by substitution method.
2x-5y=-1
x-3y=-3

Answer by ewatrrr(24785)


Hi
how do i solve this system by substitution method.
x-3y=-3 OR x = 3y - 3
2x-5y=-1
2(3y-3) - 5y = -1
y = 5 and x = 12 Ordered pair(12,5) the solution for this system